The MacWend Drive-In has determined that demand for hamburgers is given by the following equation:

Q = 205.2 + 23.0A - 200.0PM + 100.0PC + 0.5I
(1.85) (2.64) (-5.61) (2.02) (4.25)

where Q is the number of hamburgers sold per month (in 1,000s), A is the advertising expenditures during the previous month (in $1,000), PM is the price of MacWend burgers (dollars), PC is the price of hamburgers of the company's major competitor (dollars), and I is income per capita in the surrounding community (in $1,000). The t-statistics for each coefficient is shown in parentheses below each coefficient.
(A)Are the signs of the individual coefficients consistent with predictions from economic theory? Explain.
(B). If A = $5,000, PM = $1, PC = $1.20, and I = $20,000, how many hamburgers will be demanded?
(C). What is the advertising elasticity at A = $5,000?
1
Expert's answer
2018-02-21T09:00:07-0500
A) No, the signs of individual coefficients do not match the forecasts of economic theory, because some of the coefficients of t-statistics are not within the limits of this econometric model. Therefore, this model needs to improve the values of t statistics to improve the results.
B) If A = $ 5,000, PM = $ 1, PC = $ 1.20, and I = $ 20,000, then Q = 205.2 + 23.0*5 - 200.0*1 + 100.0*1.20 + 0.5*20= 250.2
C) AED = (delQ/Q)/(delA/A) = ((23/250.2) / (1/5)) = 0.46